OpenStudy (anonymous):

(CD) on a line with opposite rays. c=-2 and d= -5 1/4

OpenStudy (anonymous):

cd=d-c =-51/4+2 =-43/4

OpenStudy (anonymous):

CD = d-c =-51/4-(-2) =-51/4+2 =-43/4

OpenStudy (anonymous):

thank u two so much